identify the vertex of the graph of the function y=(x+3) square root-2.
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(28593)

The vertex of y = a(x-h)^2+k is (h,k).


In the case of y = (x+3)^2-2, we see that h = -3 and k = -2


So the vertex is (-3, -2)

Answer by stanbon(57307)

identify the vertex of the graph of the function y=(x+3) square root-2.
----
If you mean y = (x+3)^2 - 2
the vertex is (-3,-2)
==========================
Cheers,
Stan H.
=================